# Under-Canopy Navigation using Aerial Lidar Maps

### v1

@@ VIII Conclusion & Future Work
We presented a navigation system for mapping using prior overhead lidar data and planning in complex forest environments. Our proposed mapping estimates the obstructed areas on ground level using a scoring function of occupancy values obtained from airborne lidar scans and the sensor trajectory. A Monte Carlo sampling approach to incorporate sensor pose uncertainty in occupancy maps is also presented and extensively tested in simulation, showing more accurate maps than the standard occupancy mapping. We proposed two path cost functions that take into account both obstruction uncertainty and path length. Lastly, D* Lite leverages the cost functions to plan optimal paths for ground-navigation, also allowing efficient replanning. Simulations and real experiments demonstrated that our navigation approach using prior information leads to shorter paths than the naive baseline (no prior) and increases the chances of reaching the goal even in challenging forests scenarios. In future work we intend to learn the obstruction score map in a self-supervised fashion online, improving the generalisation and autonomy of our approach.

### v2

@@ VIII Conclusion & Future Work
We presented a navigation system for mapping using prior overhead lidar data and planning in complex forest environments. Our proposed mapping estimates the obstructed areas on ground level using a scoring function of occupancy values obtained from airborne lidar scans and the sensor trajectory. A Monte Carlo sampling approach to incorporate sensor pose uncertainty in occupancy maps is also presented and extensively tested in simulation, showing more accurate maps than the standard occupancy mapping. We proposed two path cost functions that take into account both obstruction uncertainty and path length. Lastly, D* Lite leverages the cost functions to plan optimal paths for ground-navigation, also allowing efficient replanning. Simulations and real experiments demonstrated that our navigation approach using prior information leads to shorter paths than the naïve baseline (no prior) and increases the chances of reaching the goal even in challenging forests scenarios. In future work we intend to learn the obstruction score map in a self-supervised fashion online and investigate more accurate trajectory sampling techniques from the posterior of modern SLAM systems (see Sec. IV-B).